OpenStudy (anonymous):

When I write an equation, I can't put large expressions in fraction. How do you write the "code" for it? I tried typing "frac" and then the expression, but that expression goes to the denominator instead to the numerator. What do you have to type then to get it right?

OpenStudy (anonymous):

Click on the left blue button "Equation", type frac{3}{4} in the window and insert . Then you will see the correct code in your answer \[\frac{3}{4}\]

OpenStudy (anonymous):

\[\alpha=\frac {\mu _{k} \times mg \times R}{\frac{2}{5} m R^2}\] Awesome!, thanks.

OpenStudy (anonymous):

You are welcome. Glad that I was able to help. I also had the same problem recently, then read somebody explanation and used it.